OpenStudy (anonymous):

Is (0,-2) a solution of 2x-5y=10?

OpenStudy (anonymous):

Well you have to plug in 0 for x and -2 for y which would be 2(0) - 5(-2) = 10 so 0+10=10 so yes 10 does equal 10. Hope you understand! And the -5* -2 = 10 because the two negatives form a positive

OpenStudy (anonymous):

Yes, 2(0)-5(-2)=10

OpenStudy (anonymous):

look : If we have : 2x-5y=10 2*0+5*(-2)=10 0+-10=10 ??? Right ?!

OpenStudy (anonymous):

but we have -5 and its right !

OpenStudy (anonymous):

The 0 cancels out the 2 leaving a -5 -5*-2 is 10

OpenStudy (anonymous):

All I needed was a "yes" or "no" answer.

OpenStudy (anonymous):

It's against the code of conduct to just give answers.

OpenStudy (anonymous):

That's true if you are taking a test. I'm not. Its just notes. And my answer choices were a "yes" or a "no"